# Efficient Methods for Non-stationary Online Learning

### v2

@@ 5.1.2 Projection Computational Complexity
=$\mathcal{M}=\Big{\{}M\triangleq(M^{[1]},\dots,M^{[H]})\in\left(\mathbb{R}^{d_{u}\times d_{x}}\right)^{H}\mid\lVert M^{[i]}\rVert_{\operatorname{op}}\leq c_{i}\Big{\}},$ (22)
@@ 5.1.3 Efficient Reduction
=$\mathcal{M}^{\prime}=\left\{M=(M^{[1]},\dots,M^{[H]})\in\left(\mathbb{R}^{d_{u}\times d_{x}}\right)^{H}\mid\lVert M^{[i]}\rVert_{\operatorname{F}}\leq c_{i}\sqrt{d}\right\},$ (23)
@@ 5.2.1 Problem Formulation
[…]
@@ 5.2.3 Efficient Reduction
10:     Meta-algorithm updates weight $\bm{p}_{t+1}\in\Delta_{\lvert A_{t+1}\rvert}$ according to Eq. (12)Eq. (13), and Eq. (14) with ^$\widehat{\ell}_{t}={\operatorname{tr}\left(\nabla g_{t}(\widehat{\mathbf{P}}^{s}_{t})\cdot\widehat{\mathbf{P}}_{t}^{s}\right)}/{2\sqrt{k}}$ and $\ell_{t,i}={\operatorname{tr}\left(\nabla g_{t}(\widehat{\mathbf{P}}^{s}_{t})\cdot\widehat{\mathbf{P}}_{t,i}^{s}\right)}/{2\sqrt{k}}$.
@@ 7 Conclusion
One important open question remains regarding another type of problem-dependent bound that scales with gradient variation, defined as $V_{T}\triangleq\sum_{t=2}^{T}\sup_{\mathbf{x}\in\mathcal{X}}\lVert\nabla f_{t}(\mathbf{x})-\nabla f_{t-1}(\mathbf{x})\rVert_{2}^{2}$ (Chiang et al., 2012; Yang et al., 2014). This bound plays a crucial role in achieving fast convergence in zero-sum games (Syrgkanis et al., 2015; Zhang et al., 2022a). While Zhao et al. (2024) have introduced a two-layer method that attains gradient-variation dynamic regret with only one gradient and one function value query per iteration, developing a projection-efficient variant in this context remains an open challenge. As mentioned, the reduction requirement is necessary but not sufficient, and integrating optimistic online learning into our reduction mechanism is notably difficult due to the constrained feasible domain and the complexity of the two-layer structure. Another important problem is to understand the minimal computational overhead required for robustness to non-stationarity. For instance, is it possible to design a single-layer algorithm that achieves optimal dynamic regret, adaptive regret, or even interval regret, similar to the ensemble methods presented in this paper? If not, information-theoretic arguments demonstrating the necessity of more complex structures would be highly interesting.
@@ A.1 Properties of Distance Function
Let $\mathcal{X}\subseteq\mathbb{R}^{d}$ a closed convex set. Given $\mathbf{y}\in\mathbb{R}^{d}$ and $\mathbf{y}\notin\mathcal{X}$. Let $\mathbf{x}=\Pi_{\mathcal{X}}[\mathbf{y}]$. Then we have $\{\frac{\mathbf{y}-\mathbf{x}}{\lVert\mathbf{y}-\mathbf{x}\rVert_{2}}\}=\partial S_{\mathcal{X}}(\mathbf{y})$.
@@ A.2 Proof of Theorem 1
Let us focus on the case when $\langle\nabla f_{t}(\mathbf{x}_{t}),\mathbf{v}_{t}\rangle<0$. First, it can be verified that the surrogate loss $g_{t}(\mathbf{y})=\langle\nabla f_{t}(\mathbf{x}_{t}),\mathbf{y}\rangle-\langle\nabla f_{t}(\mathbf{x}_{t}),\mathbf{v}_{t}\rangle\cdot S_{\mathcal{X}}(\mathbf{y})$ is convex due to the convexity of $S_{\mathcal{X}}(\mathbf{y})$ shown in Lemma 2 and the condition of $\langle\nabla f_{t}(\mathbf{x}_{t}),\mathbf{v}_{t}\rangle<0$ in this case. Next, the gradient of $g_{t}(\cdot)$ at the $\mathbf{y}_{t}$ point can be calculated according to Lemma 1 as,
@@ B.1 Proof of Theorem 2
Notice that Zhang et al. (2018a) propose the improved Ader algorithm (see Algorithm 3 and Algorithm 4 in their paper), which uses the linearized loss as the input to make the online algorithm requiring one gradient and one function evaluation per iteration. So the algorithm satisfies the requirements of our reduction mechanism, and our algorithm (Algorithm 1 with specifications in Theorem 1) can be regarded as the improved Ader equipped with the projection-efficient reduction. As a consequence, we can directly obtain the same dynamic regret guarantee and meanwhile ensure $1$ projection complexity by following the same proof of the improved Ader as well as the reduction guarantee (Theorem 1). ∎
@@ B.2 Proof of Theorem 3
Next, we specify the base-learner $\mathcal{E}_{i}$ compared with. Indeed, we aim at choosing the one with step size closest to the (near-)optimal step size $\eta^{\star}=\sqrt{\frac{5D^{2}+2DP_{T}}{1+8LF_{T}^{\mathbf{x}}}}$, where we denote by $F_{T}^{\mathbf{x}}=\sum_{t=1}^{T}f_{t}(\mathbf{x}_{t})$ the cumulative loss of the decisions. By Assumption 1 and Assumption 2, we have $F_{T}^{\mathbf{x}}\in[0,GDT]$ and then the possible minimum optimal and maximum step size are
@@
The construction of step size pool is by discretizing the interval $[\eta_{\min},\eta_{\max}]$ with intervals with exponentially increasing length. The step size of each base-learner is designed to be monotonically increasing with respect to the index. Consequently, it is evident to verify that there exists an index $i^{\star}\in[N]$ such that $\eta_{i^{\star}}\leq\eta^{\star}\leq\eta_{i^{\star}+1}=2\eta_{i^{\star}}$. As the upper bounds of meta-regret and base-regret hold for any compared base-learner, we can choose the index as $i^{\star}$ in particular. Then the second and the third terms in the inequality (32) satisfy
@@ C.1 Key Lemmas
$\leq 4D\bigg{(}3\sqrt{\ln(1+2m)}+\frac{\mu_{t}}{\sqrt{\ln(1+2m)}}\bigg{)}\sqrt{L\sum_{\tau=i}^{t}f_{t}(\mathbf{x}_{t})}+18GD\ln(1+2m)+6GD\mu_{t}$
@@
Under Assumptions 12, and 3, for any interval $[i,j]\in\widetilde{\mathcal{C}}$ in the geometric covering intervals defined in Eq. (11), at the beginning of which we assume the $m$-th base-learner is initialized, Algorithm 2 ensures for any time $t\in[i,j]$ and any comparator $\mathbf{u}\in\mathcal{X}$,
@@
It is worth emphasizing that, the regret bound in terms of the cumulative loss of final decision $\{\mathbf{x}_{t}\}$ plays a key role in proving the worst-case adaptive regret bound in Theorem 4.
@@
Under Assumptions 12, and 3, for any interval $[s_{m},s_{m+1}-1]$ determined by two consecutive intervals $s_{m}$ and $s_{m+1}$, where we denote by $s_{m}$ the $m$-th marker, Algorithm 2 ensures that for any comparator $\mathbf{u}\in\mathcal{X}$,
@@
Under Assumptions 12, and 3, for any interval $[i,j]\in\widetilde{\mathcal{C}}$ and any $t\in[i,j]$, the variable $m$ specified in Lemma 4 and Lemma 5 can be bounded by
@@
$\mu_{t}=\ln(1+(1+\ln(1+t))/(2e))$, and $F_{[a,b]}=\mathop{\min}_{\mathbf{u}\in\mathcal{X}}\sum_{\tau=a}^{b}f_{\tau}(\mathbf{u})$ denotes the cumulative loss of the comparator within the interval $[a,b]\subseteq[T]$.
@@ C.2 Proof of Theorem 4
Given any interval $[r,s]\subseteq[T]$, we will identify a series of intervals $I_{1},\dots,I_{v}$ in the schedule $\widetilde{\mathcal{C}}$ that almost covers the entire interval $[r,s]$. Then, we can use Lemma 7 to ensure low regret over these intervals. By further demonstrating that the regret on the uncovered interval can be well-controlled and that the number of intervals $v$ is not large, we can ultimately achieve the desired bound. Below, we provide formal proof.
Recall in the algorithmic procedures, the algorithm will register a series markers $s_{1},s_{2},\ldots$. Let $s_{p}$ be the smallest marker that is larger than $r$, and let $s_{q}$ be the largest marker that is not large than $s$. As a result, we have $s_{p-1}\leq r<s_{p}$, and $s_{q}\leq s<s_{q+1}$.
We bound the regret over the interval $[r,s_{p}-1]$ that is not covered by the schedule as
@@
Next, we focus on the interval $[s_{p},s]$, which can be covered by geometric covering intervals. By Lemma 19, we can find $v$ consecutive intervals
@@
Through $[s_{p},s_{q}-1]$, the algorithm registers $q-p$ markers, i.e., $s_{p},s_{p+1},\dots,s_{q-1}$, then by Lemma 6 we can lower bound the cumulative loss of comparator $\mathbf{u}^{\star}\triangleq\operatorname*{arg\,min}_{\mathbf{u}\in\mathcal{X}}\sum_{t=r}^{s}f_{t}(\mathbf{u})$ by the corresponding thresholds,
@@
where we use the notation $F^{\mathbf{x}}_{[a,b]}=\sum_{t=a}^{b}f_{t}(\mathbf{x}_{t})$ to denote the cumulative loss of the final decisions within the interval $[a,b]\subseteq[T]$, and we apply Lemma 7 to upper bound $m\leq\mathcal{O}(T)$ as only the worst-case behavior matters now.
@@
$\sum_{t=s_{i_{k}}}^{s_{i_{k+1}}-1}f_{t}(\mathbf{x}_{t})$ $=\sum_{a=i_{k}}^{i_{k+1}-1}\left(\Bigg{(}\sum_{t\in[s_{p},s_{p+1}-2]}f_{t}(\mathbf{x}_{t})\Bigg{)}+f_{s_{a+1}-1}(\mathbf{x}_{s_{a+1}-1})\right)$
@@
We then split a given interval $[r,s]$ into three parts to analyze, namely, the consecutive $v-1$ intervals $I_{1}$ to $I_{v-1}$, interval $[r,s_{p}-1]$, and $[s_{i_{v}},s]$, where notably the last two intervals are not fully covered by any interval in geometric covering intervals,
@@
$\leq\mathcal{O}\left(v\log T+\sqrt{\log T\cdot C_{i_{v}-1}\cdot\lvert i_{v}-i_{v-1}\rvert}\right),$ (46)
@@
In the subsequent analysis, our objective is to demonstrate that $C_{i_{v}-1}\cdot|i_{v}-i_{v-1}|=\mathcal{O}(\lvert I\rvert)$. Employing the mechanism of the time-varying threshold as defined in Eq. (15), we have
@@
[…]
@@ C.3 Proof of Lemma 4
First we introduce some useful variables to help us prove the adaptivity of Adapt-ML-Prod under sleeping-expert setting. Similar to the proof technique of Daniely et al. (2015), for any interval $[i,j]\in\widetilde{\mathcal{C}}$ in the geometric covers defined in (11), on which we suppose the $m$-th base-learner is active, we define the following pseudo-weight for the $m$-th base-learner,
$\widetilde{w}_{\tau,m}=\left\{\begin{array}[]{cl}0&\tau<i,\\ 1&\tau=i,\\ \big{(}\widetilde{w}_{\tau-1,m}(1+\eta_{\tau-1}(\widehat{\ell}_{\tau-1}-\ell_{\tau-1,m}))\big{)}^{\frac{\eta_{\tau,m}}{\eta_{\tau-1,m}}}&i<\tau\leq j+1,\\ \widetilde{w}_{j+1,m}&\tau>j+1.\end{array}\right.$
In addition, we use $\widetilde{W}_{t}=\sum_{k\in[T]}\widetilde{w}_{t,k}$ to denote the summation of pseudo-weights for all possible base-learners up to time $t$. As for the problem-dependent geometric covers, in the worst case there are at most $T$ base-learners generated, we use $[T]$ to denote the indexes for all the base-learners. Notice that the pseudo-weight $\widetilde{w}_{t}$ is defined as 00 for asleep base-learners till time $t$, so we can include all possible ones safely in the definition even though they are not generated in practical implementations of the algorithm.
@@
By the construction of the geometric covers as specified in Eq. (11), we know that there will be at most $2m$ base-learners initialized for the $m$-th base-learner active on interval $[i,j]$ till her end. This is because $m$-th base-learner is initialized when $m$-th marker is recorded, and she will expire before the moment when $2m$-th marker is recorded, as demonstrated by the construction of cover defined in Eq. (11). Owing to this property, we have $\widetilde{W}_{t+1}=\sum_{k\in[2m]}\widetilde{w}_{t+1,k}$ as others’ pseudo-weight equals to 00 by definition. So we can upper bound $\widetilde{W}_{t+1}$ as,
@@
where we denote by $[i_{k},j_{k}]\in\widetilde{\mathcal{C}}$ the active time for $k$-th base-learner.
@@
where the first equality holds by the definition of $\widetilde{w}_{t+1,k}$, the second inequality is by the updating rule of $\widetilde{w}_{t+1,k}$ and Lemma (26), and the second term in the last equality equals to 00 due to the weight update rule in (14) and the fact of $\widetilde{w}_{t,k}=w_{t,k}$ for any $t\in[i_{k},j_{k}]$.
@@
$\sum_{\tau=i}^{t}\langle\nabla g_{t}(\mathbf{y}_{t}),\mathbf{y}_{t}-\mathbf{y}_{t,m}\rangle$

### v3

@@ 5.1.2 Projection Computational Complexity
$\mathcal{M}=\Big{\{}M\triangleq(M^{[1]},\dots,M^{[H]})\in\left(\mathbb{R}^{d_{u}\times d_{x}}\right)^{H}\mid\lVert M^{[i]}\rVert_{\operatorname{op}}\leq c_{i}\Big{\}},$ (22)
@@ 5.1.3 Efficient Reduction
$\mathcal{M}^{\prime}=\left\{M=(M^{[1]},\dots,M^{[H]})\in\left(\mathbb{R}^{d_{u}\times d_{x}}\right)^{H}\mid\lVert M^{[i]}\rVert_{\operatorname{F}}\leq c_{i}\sqrt{d}\right\},$ (23)
@@ 5.2.1 Problem Formulation
$\textsc{A-Reg}(|I|)=\max_{[r,r+\tau-1]\subseteq[T]}\left\{\sum_{t=r}^{r+\tau-1}f_{t}(\mathbf{P}_{t})-\min_{\mathbf{P}\in\mathcal{P}_{k}}\sum_{t=r}^{r+\tau-1}f_{t}(\mathbf{P})\right\}.$ (26)
Yuan and Lamperski (2019) examine a regret notion similar to but weaker than Eq. (26), defined as $\textsc{WA-Reg}(T)=\max_{[r,q]\subseteq[T]}\left\{\sum_{t=r}^{q}f_{t}(\mathbf{P}_{t})-\min_{\mathbf{P}\in\mathcal{P}_{k}}\sum_{t=r}^{q}f_{t}(\mathbf{P})\right\}$. This variant is usually termed as the weakly adaptive regret (Hazan and Seshadhri, 2007; Zhang et al., 2018b), which lacks the guarantee for intervals with $|I|\leq\mathcal{O}(\sqrt{T})$. It should be noted that Yuan and Lamperski (2019) propose an algorithm with an $\widetilde{\mathcal{O}}(\sqrt{T})$ weakly adaptive regret for online PCA, and to the best of our knowledge, current literature lacks algorithms with an $\widetilde{\mathcal{O}}(\sqrt{|I|})$ strongly adaptive regret for online PCA. We not only design the first algorithm with such a strongly adaptive regret guarantee, but also implement our reduction mechanism to ensure that it enjoys a projection complexity of $1$.
@@ 5.2.3 Efficient Reduction
10:  Meta-algorithm updates weight $\bm{p}_{t+1}\in\Delta_{\lvert A_{t+1}\rvert}$ according to Eq. (12)Eq. (13), and Eq. (14) with $\widehat{\ell}_{t}={\operatorname{tr}\left(\nabla g_{t}(\widehat{\mathbf{P}}^{s}_{t})\cdot\widehat{\mathbf{P}}_{t}^{s}\right)}/{2\sqrt{k}}$ and $\ell_{t,i}={\operatorname{tr}\left(\nabla g_{t}(\widehat{\mathbf{P}}^{s}_{t})\cdot\widehat{\mathbf{P}}_{t,i}^{s}\right)}/{2\sqrt{k}}$.
@@ 7 Conclusion
One important open question remains regarding another type of problem-dependent bound that scales with gradient variation, defined as $V_{T}\triangleq\sum_{t=2}^{T}\sup_{\mathbf{x}\in\mathcal{X}}\lVert\nabla f_{t}(\mathbf{x})-\nabla f_{t-1}(\mathbf{x})\rVert_{2}^{2}$ (Chiang et al., 2012; Yang et al., 2014). This bound plays a crucial role in achieving fast convergence in zero-sum games (Syrgkanis et al., 2015; Zhang et al., 2022a). While Zhao et al. (2024) introduced a two-layer method that attains gradient-variation dynamic regret with only one gradient and one function value query per iteration, developing a projection-efficient variant in this context remains an open challenge. As mentioned, the reduction requirement is necessary but not sufficient, and integrating optimistic online learning into our reduction mechanism is notably difficult due to the constrained feasible domain and the complexity of the two-layer structure. Another important problem is to understand the minimal computational overhead required for robustness to non-stationarity. For instance, is it possible to design a single-layer algorithm that achieves optimal dynamic regret, adaptive regret, or even interval dynamic regret, similar to the ensemble methods presented in this paper? Preliminary results suggest that a single-layer algorithm with a wavelet-based restarting scheme can achieve optimal dynamic regret in certain regimes, provided the online learner has access to noisy feedback about the comparators (Qian et al., 2024). However, its applicability to more general settings remains unclear. Extending this approach to broader scenarios would be valuable. Alternatively, information-theoretic arguments demonstrating the necessity of more complex structures would also be highly interesting.
Acknowledgment
This work was supported by National Science and Technology Major Project (2022ZD0114800) and NSFC (U23A20382). Peng Zhao was supported in part by the Xiaomi Foundation. We are grateful for anonymous reviewers for their careful reviews and many constructive suggestions.
@@ A.1 Properties of Distance Function
Let $\mathcal{X}\subseteq\mathbb{R}^{d}$ be a closed convex set. Given $\mathbf{y}\in\mathbb{R}^{d}$ and $\mathbf{y}\notin\mathcal{X}$. Let $\mathbf{x}=\Pi_{\mathcal{X}}[\mathbf{y}]$. Then we have $\{\frac{\mathbf{y}-\mathbf{x}}{\lVert\mathbf{y}-\mathbf{x}\rVert_{2}}\}=\partial S_{\mathcal{X}}(\mathbf{y})$.
@@ A.2 Proof of Theorem 1
Let us focus on the case when $\langle\nabla f_{t}(\mathbf{x}_{t}),\mathbf{v}_{t}\rangle<0$. First, it can be verified that the surrogate loss $g_{t}(\mathbf{y})=\langle\nabla f_{t}(\mathbf{x}_{t}),\mathbf{y}\rangle-\langle\nabla f_{t}(\mathbf{x}_{t}),\mathbf{v}_{t}\rangle\cdot S_{\mathcal{X}}(\mathbf{y})$ is convex due to the convexity of $S_{\mathcal{X}}(\mathbf{y})$ shown in Lemma 2 and the condition of $\langle\nabla f_{t}(\mathbf{x}_{t}),\mathbf{v}_{t}\rangle<0$ in this case. Next, the gradient of $g_{t}(\cdot)$ at the $\mathbf{y}_{t}$ point can be calculated according to Lemma 1 as
@@ B.1 Proof of Theorem 2
Notice that Zhang et al. (2018a) propose the improved Ader algorithm (see Algorithm 3 and Algorithm 4 in their paper), which uses the linearized loss as the input which allows the online algorithm to require one gradient and one function value. So the algorithm satisfies the requirements of our reduction mechanism, and our algorithm (Algorithm 1 with specifications in Theorem 1) can be regarded as the improved Ader equipped with the projection-efficient reduction. As a consequence, we can directly obtain the same dynamic regret guarantee and meanwhile ensure $1$ projection complexity by following the same proof of the improved Ader as well as the reduction guarantee (Theorem 1). ∎
@@ B.2 Proof of Theorem 3
Next, we specify the base-learner $\mathcal{E}_{i}$ to be compared with. Indeed, we aim at choosing the one with step size closest to the (near-)optimal step size $\eta^{\star}=\sqrt{\frac{5D^{2}+2DP_{T}}{1+8LF_{T}^{\mathbf{x}}}}$, where we denote by $F_{T}^{\mathbf{x}}=\sum_{t=1}^{T}f_{t}(\mathbf{x}_{t})$ the cumulative loss of the decisions. By Assumption 1 and Assumption 2, we have $F_{T}^{\mathbf{x}}\in[0,GDT]$ and then the possible minimum optimal and maximum step size are
@@
The construction of step size pool is by discretizing the interval $\eta_{\min},\eta_{\max}]$ with intervals with exponentially increasing length. The step size of each base-learner is designed to be monotonically increasing with respect to the index. Consequently, it is evident to verify that there exists an index $i^{\star}\in[N]$ such that $\eta_{i^{\star}}\leq\eta^{\star}\leq\eta_{i^{\star}+1}=2\eta_{i^{\star}}$. As the upper bounds of meta-regret and base-regret hold for any compared base-learner, we can choose the index as $i^{\star}$ in particular. Then the second and the third terms in the inequality (32) satisfy
@@ C.1 Key Lemmas
$\leq 4D\bigg{(}3\sqrt{\ln(1+2m)}+\frac{\mu_{t}}{\sqrt{\ln(1+2m)}}\bigg{)}\sqrt{L\sum_{\tau=i}^{t}f_{\tau}(\mathbf{x}_{\tau})}+18GD\ln(1+2m)+6GD\mu_{t}$
@@
Under Assumptions 12, and 3, for any interval $i,j]\in\widetilde{\mathcal{C}}$ in the geometric covering intervals defined in Eq. (11), at the beginning of which we assume the $m$-th base-learner is initialized, Algorithm 2 ensures for any time $t\in[i,j]$ and any comparator $\mathbf{u}\in\mathcal{X}$,
@@
It is worth emphasizing that, the regret bound in terms of the cumulative loss of final decisions $\{\mathbf{x}_{t}\}$ plays a key role in proving the worst-case adaptive regret bound in Theorem 4.
@@
Under Assumptions 12, and 3, for any interval $s_{m},s_{m+1}-1]$ determined by two consecutive intervals $s_{m}$ and $s_{m+1}$, where we denote by $s_{m}$ the $m$-th marker, Algorithm 2 ensures that for any comparator $\mathbf{u}\in\mathcal{X}$,
@@
Under Assumptions 12, and 3, for any interval $i,j]\in\widetilde{\mathcal{C}}$ and any $t\in[i,j]$, the variable $m$ specified in Lemma 4 and Lemma 5 can be bounded by
@@
$\mu_{t}=\ln(1+(1+\ln(1+t))/(2e))$, and $F_{[a,b]}=\mathop{\min}_{\mathbf{u}\in\mathcal{X}}\sum_{\tau=a}^{b}f_{\tau}(\mathbf{u})$ denotes the cumulative loss of the comparator within the interval $a,b]\subseteq[T]$.
@@ C.2 Proof of Theorem 4
Given any interval $r,s]\subseteq[T]$, we will identify a series of intervals $I_{1},\dots,I_{v}$ in the schedule $\widetilde{\mathcal{C}}$ that almost covers the entire interval $[r,s]$. Then, we can use Lemma 7 to ensure low regret over these intervals. By further demonstrating that the regret on the uncovered interval can be well-controlled and that the number of intervals $v$ is not large, we can ultimately achieve the desired bound. Below, we provide formal proof.
Recall in the algorithmic procedures, the algorithm will register a series of markers $s_{1},s_{2},\ldots$. Let $s_{p}$ be the smallest marker that is larger than $r$, and let $s_{q}$ be the largest marker that is not larger than $s$. As a result, we have $s_{p-1}\leq r<s_{p}$, and $s_{q}\leq s<s_{q+1}$.
We bound the regret over the interval $r,s_{p}-1]$ that is not covered by the schedule as
@@
Next, we focus on the interval $s_{p},s]$, which can be covered by geometric covering intervals. By Lemma 19, we can find $v$ consecutive intervals
@@
Through $s_{p},s_{q}-1]$, the algorithm registers qpq-p markers, i.e., $s_{p},s_{p+1},\dots,s_{q-1}$, then by Lemma 6 we can lower bound the cumulative loss of comparator $\mathbf{u}^{\star}\triangleq\operatorname*{arg\,min}_{\mathbf{u}\in\mathcal{X}}\sum_{t=r}^{s}f_{t}(\mathbf{u})$ by the corresponding thresholds,
@@
where we use the notation $F^{\mathbf{x}}_{[a,b]}=\sum_{t=a}^{b}f_{t}(\mathbf{x}_{t})$ to denote the cumulative loss of the final decisions within the interval $a,b]\subseteq[T]$, and we apply Lemma 7 to upper bound $m\leq\mathcal{O}(T)$ as only the worst-case behavior matters now.
@@
$\sum_{t=s_{i_{k}}}^{s_{i_{k+1}}-1}f_{t}(\mathbf{x}_{t})$ $=\sum_{a=i_{k}}^{i_{k+1}-1}\left(\Bigg{(}\sum_{t\in[s_{a},s_{a+1}-2]}f_{t}(\mathbf{x}_{t})\Bigg{)}+f_{s_{a+1}-1}(\mathbf{x}_{s_{a+1}-1})\right)$
@@
We then split a given interval $[r,s]$ into three parts to analyze, namely, the consecutive v1v-1 intervals $I_{1}$ to $I_{v-1}$, interval $r,s_{p}-1]$, and $s_{i_{v}},s]$, where notably the last two intervals are not fully covered by any interval in geometric covering intervals,
@@
[…]
@@
[…]
@@
$\texttt{term-(c)}\leq\mathcal{O}\left(\log T+\sqrt{F_{[s_{i_{v}},s]}\log T}\right)\leq\mathcal{O}\left(\log T+\sqrt{|I|\log T}\right).$
@@ C.3 Proof of Lemma 4
First we introduce some useful variables to help us prove the adaptivity of Adapt-ML-Prod under sleeping-expert setting. Similar to the proof technique of Daniely et al. (2015), for any interval $i,j]\in\widetilde{\mathcal{C}}$ in the geometric covers defined in (11), on which we suppose the $m$-th base-learner is active, we define the following pseudo-weight for the $m$-th base-learner,
$\widetilde{w}_{\tau,m}=\left\{\begin{array}[]{cl}0&\tau<i,\\ 1&\tau=i,\\ \big{(}\widetilde{w}_{\tau-1,m}(1+\eta_{\tau-1,m}(\widehat{\ell}_{\tau-1}-\ell_{\tau-1,m}))\big{)}^{\frac{\eta_{\tau,m}}{\eta_{\tau-1,m}}}&i<\tau\leq j+1,\\ \widetilde{w}_{j+1,m}&\tau>j+1.\end{array}\right$.
In addition, we use $\widetilde{W}_{t}=\sum_{k\in[T]}\widetilde{w}_{t,k}$ to denote the summation of pseudo-weights for all possible base-learners up to time $t$. As for the problem-dependent geometric covers, in the worst case there are at most $T$ base-learners generated, we use $[T]$ to denote the indexes for all the base-learners. Notice that the pseudo-weight $\widetilde{w}_{t}$ is defined as 0 for asleep base-learners till time $t$, so we can include all possible ones safely in the definition even though they are not generated in practical implementations of the algorithm.
@@
By the construction of the geometric covers as specified in Eq. (11), we know that there will be at most $2m$ base-learners initialized for the $m$-th base-learner active on interval $[i,j]$ till her end. This is because $m$-th base-learner is initialized when $m$-th marker is recorded, and she will expire before the moment when $2m$-th marker is recorded, as demonstrated by the construction of cover defined in Eq. (11). Owing to this property, we have $\widetilde{W}_{t+1}=\sum_{k\in[2m]}\widetilde{w}_{t+1,k}$ as others’ pseudo-weight equals to 0 by definition. So we can upper bound $\widetilde{W}_{t+1}$ as,
@@
where we denote by $i_{k},j_{k}]\in\widetilde{\mathcal{C}}$ the active time for $k$-th base-learner.
@@
where the first equality holds by the definition of $\widetilde{w}_{t+1,k}$, the second inequality is by the updating rule of $\widetilde{w}_{t+1,k}$ and Lemma (26), and the second term in the last equality equals to 0 due to the weight update rule in (14) and the fact of $\widetilde{w}_{t,k}=w_{t,k}$ for any $t\in[i_{k},j_{k}]$.
@@
$\sum_{\tau=i}^{t}\langle\nabla g_{t}(\mathbf{y}_{\tau}),\mathbf{y}_{\tau}-\mathbf{y}_{\tau,m}\rangle$